OpenStudy (anonymous):

Solve the equation: 4x2 = 9 Answer {-2, -3} {2/3, -2/3} {3/2, -3/2} no solution

OpenStudy (anonymous):

@E.ali could you please?? :)))

OpenStudy (anonymous):

What is the question ????

OpenStudy (anonymous):

Solve the equation: 4x2 = 9

OpenStudy (anonymous):

amm ... 4x . 2 = 9 or 4x+2=9 ???? your answers and question ...are not same ..???

OpenStudy (anonymous):

haha

OpenStudy (anonymous):

this is question: Solve the equation: 4x2 = 9

OpenStudy (anonymous):

these are the answers: {-2, -3} {2/3, -2/3} {3/2, -3/2} no solution

OpenStudy (anonymous):

..??

OpenStudy (anonymous):

I think its wrong :(

OpenStudy (anonymous):

for sure not :D

OpenStudy (anonymous):

maybe no solution?

OpenStudy (unklerhaukus):

\[4x^2 = 9\] divide both sides by four, then take the square root

OpenStudy (anonymous):

@UnkleRhaukus:I think your right :)

OpenStudy (unklerhaukus):

when \(x^2\) is copy pasted it becomes x2, to be more clear it is better to write x^2, or x^(2) ,

OpenStudy (anonymous):

Yea :)

OpenStudy (anonymous):

:O

OpenStudy (anonymous):

so its no solution?

OpenStudy (unklerhaukus):

4x^2 =9 what do you get after dividing both sides by four ?

OpenStudy (unklerhaukus):

... 4x^2/4 =9/4 ...